there are 14 blue and 28 red bulbs to be used for a birthday party. they are to be placed in a plastic bag so that each bag contains the same number. what is the largest plastic bag will be needed


1
Expert's answer
2021-10-12T10:26:09-0400

Solution:

No. of blue bulbs = 14

No. of red bulbs = 28

Largest no. of plastic bag required is HCF of 14, 28.

"14=2\times 7
\\28=2\times 2\times 7"

HCF(14, 28)"=2\times 7=14" .